Solve the following measurements:

a. On a certain map, the straight-line distance between The Pas and Winnipeg is 4.1cm. If the scale of the map is 1cm = 120km, how far apart are The Pas and Winnipeg, to the nearest kilometre?


b. On a scale drawing of a house, the scale is 1cm: 4m. If the actual width of the house is 30m, what is the length of the line in the drawing that represents the width of the house?


c. If an electronics company wants to shrink the dimensions of a Tablet measuring 20cm wide by 25cm high by a scale factor of 2/5, what is the new width of the Tablet? (Recall that the Scale Factor is given by either model/original OR model:original.)


(Need to show work)


1
Expert's answer
2022-01-31T16:22:50-0500

a. Since the scale of the map is 1cm = 120km, and the straight-line distance between The Pas and Winnipeg on this map is 4.1cm, the real distance is equal to "4.1\times 120=492" km.


b. The length of the line in the drawing that represents the width of the house is

"30\/4=7.5" cm.


c. After shrinking, the new dimensions of a Tablet will be

"2/5\times 20=8" cm wide by "2/5\times 25=10" cm high.